Find the (a) mean, (b) median, (c) mode, and (d) midrange for the data and then (e) answer the given question.

Listed below are the jersey numbers of 11 players randomly selected from the roster of a championship sports team. What do the results tell us? 

9 94 58 38 42 73 46 15 88 64 7


a. Find the mean. The mean is _______ . (Type an integer or a decimal rounded to one decimal place as needed.) 

b. Find the median. The median is _______ . (Type an integer or a decimal rounded to one decimal place as needed.) 

c. Find the mode.

Select the correct choice below and, if necessary, fill in the answer box to complete your choice.

A. The mode(s) is(are)

B. There is no mode.

d. Find the midrange.

The midrange is

(Type an integer or a decimal rounded to one decimal place as needed.)

e. What do the results tell us?

A. The midrange gives the average (or typical) jersey number, while the mean and median give two different interpretations of the spread of possible jersey numbers.

B. The jersey numbers are nominal data and they do not measure or count anything, so the resulting statistics are meaningless.

C. Since only 11 of the jersey numbers were in the sample, the statistics cannot give any meaningful results.

D. The mean and median give two different interpretations of the average (or typical) jersey number, while the midrange shows the spread of possible jersey numbers.
Answer #1

Data

9,94,58,38,42,73,46,15,88,64,7

Mean = sum / n = 534 / 11 = 48.5

Arrange the data

94, 88, 73, 64, 58, 46, 42, 38, 15, 9, 7

median = middle most value = 46

Mode = there is no mode

All values appears just once

Midrange = ( low + high )/2 = ( 7+94 ) / 2 = 50.5

Answer :- D